Monochromatic light (l = 500 nm) is incident on a soap bubble (n = 1.40). What is the wavelength of the light (in nm) in the bubble film?


1
Expert's answer
2021-08-29T16:55:37-0400

The refractive index of air is 1.

The refractive index of the bubble is 1.4.

The relation:

"n_1λ_1=n_2λ_2 \\

1 \times 500 \;nm = 1.4 \times λ_2 \\

λ_2 = \frac{500}{1.4}= 357.1 \;nm"

Answer: 357.1 nm